# Efficient Multi-Agent Collaboration with Tool Use for Online Planning in Complex Table Question Answering

### v1

@@ 6 Analysis
We explore the effectiveness of specialized agents and tool use in MACT by conducting an ablation study with three scenarios: ablating only $T_{srh}$ (Wikipedia search API), ablating $T_{srh}$ and $T_{cal}$ (calculator), and further ablating the coding agent $M_{c}$ with a Python interpreter. In cases where $M_{c}$ or/and tools are ablated, the most frequent estimated observations from $M_{p}$ are used as the final observations. Our results in Table 6 show that both the tools and the coding agent contribute to the performance of the framework. Nevertheless, they contribute differently to the final performance. For instance, ablating the search tool barely influences the results whereas there are large performance drops when further ablating the coding agent and the Python interpreter. This might be due to the frequency of the usage of tools and the coding agent. We find that the search tool is barely used whereas the coding agent is called in almost every query. Detailed analysis can be found in A.5. In addition, since Wikipedia is a common pre-trained corpus for LLMs, most information might have already been encoded. Nevertheless, the search tool can still be helpful given LLMs are known to suffer from hallucinations and the knowledge encoded might not be updated in time. We further observe that the ablation affects WTQ and TAT more than CRT and SCITAB. This might be attributed to dataset features: CRT contains many yes-no questions and SCITAB has been converted from a ternary classification dataset. Thus, chances for guessing the correct final answers are higher than in datasets with a more diverse answer distribution, such as WTQ and TAT. By evaluating our framework on instances from CRT that have answers other than yes/no, we find a performance drop of 8.23 when ablating both tools and the coding agent.
@@
Figure 2: EM (bar chart) and iteration ratio (line chart) against different $\alpha$. The iteration ratio is calculated by dividing the number of iterations when using efficiency optimization with a specific $\alpha$ by the number of iterations when not using it.
To investigate the trade-off between efficiency and accuracy, we plot the model performance against $\alpha\in\{0.2,0.4,0.6,0.8,1\}$. We also plot the ratio of the total number of iterations taken to terminate from each tested $\alpha$ to the total number of iterations taken when not using the optimization, i.e., letting the planning agent decide via the Finish action when to stop the execution. Figure 2 shows that the best performance is reached for $\alpha=1$, i.e., when requiring all estimated results to agree with each other in order to stop the iteration. For SCITAB, for instance, we save approximately 40% of the iterations when setting $\alpha$ to 1 without losing performance compared to not using the optimization component (59.8 vs. 59.7) On average, adding the efficiency optimization module saves up to 33% of iterations. This shows the effectiveness of the optimization in general and that users can individually tune the desired trade-off of performance and computation time.
Error Analysis
We randomly sample 50 instances that MACT fails per dataset and conduct an error analysis. About half of the errors come from the invalid or wrong code generated by the coding agent $M_{c}$. Either $M_{c}$ fails to understand instructions, thus generating wrong code, or due to the complex table data type, the code execution is not successful. This indicates the importance of table pre-processing. The second error type can be attributed to evaluation. We find that about one-third of failures come from strict evaluation metrics (Exact Match Accuracy). This influences the performance of MACT on the TAT dataset the most, as it features long text strings as answers. The evaluation challenge has been discussed in many previous works Wu and Feng (2024); Li et al. (2024). To estimate the upper-bound of our method, we use GPT-4 for evaluation and achieve an accuracy of 87.8% on the TAT dataset (66.2% with EM). The remaining error cases can be largely attributed to the failure of the planning agent. This means the planning agent does not decompose questions correctly. For instance, one question asks the score range (min-max) for the top 10 finishers. Apart from retrieving the min and max scores of the top 10 finishers, the planner continues to generate the action: Calculate[Calculate the range of the scores in the observation 1.]. This results in the wrong final prediction of the difference between the min and max scores.
@@ 7 Conclusions
We propose multi-agent collaboration with tool use for table question answering. Unlike previous work, our MACT framework neither requires fine-tuning nor does it depend on closed-source models. In our experiments, our framework demonstrates generalizability across different benchmark datasets and outperforms a number of state-of-the-art approaches, including closed-source commercial models and fine-tuned models. To boost efficiency, we introduce an efficiency optimization module that saves up to 33% of the iterations in our analysis. In our experiments and analyses, we show that multi-agent collaboration with tools is an effective approach for table question answering.
8 Limitation
@@ A.1 Datasets
Table A.1 shows their statistics and characteristics. WTQ Pasupat and Liang (2015), TAT (Zhu et al., 2021), CRT Zhang et al. (2023b) and SCITAB (Lu et al., 2023) are publicly available under the licenses of CC-BY-SA-4.05, MIT and MIT, respectively. These licenses all permit us to compose, modify, publish, and dis- tribute additional annotations upon the original dataset.
@@ A.3 Analysis of Iteration Number Distribution
A.5 Analysis of Intent Distribution
We report the distribution of the intents of the selected actions for each dataset in Figure 3. We observe that Retrieve and Calculate are the most frequent intents, along with Finish. This indicates that our proposed second agent, the coding agent $M_{c}$ is used frequently. Different datasets also require different intents. In particular, the framework needs to use the intent Read for solving instances in TAT, where textual descriptions are given while this is not the case for the other datasets. We notice that the Search intent is used only few times across datasets. This might be because most instances were designed to be solved using the given table and text information. However, when looking into the individual cases where Search is used, we still find it useful. For instance, one question from WTQ asks about the number of athletes from American but no information about nationality is given in the table. In this case, Search assists with answering the question by adding the nationality for each athlete from Wikipedia.
Figure 3: The distribution of action intents for different datasets.
@@ A.6 MACT Prompts
We provide the prompts used for the planning agent for the examined datasets (WTQ, TAT, CRT and SCITAB) in Figures 6, 7, 8, and 9. Figure 10 and Figure 11 show the prompts used for the coding agent for the action intents Retrieval and Calculation, respectively.

### v2

@@ 6 Analysis
We explore the effectiveness of specialized agents and tool use in MACT by conducting an ablation study with three scenarios: ablating only $T_{srh}$ (Wikipedia search API), ablating $T_{srh}$ and $T_{cal}$ (calculator), and further ablating the coding agent $M_{c}$ with a Python interpreter. In cases where $M_{c}$ or/and tools are ablated, the most frequent estimated observations from $M_{p}$ are used as the final observations. Our results in Table 6 show that both the tools and the coding agent contribute to the performance of the framework. Nevertheless, they contribute differently to the final performance. For instance, ablating the search tool barely influences the results whereas there are large performance drops when further ablating the coding agent and the Python interpreter. We find that the search tool is barely used whereas the coding agent is called in almost every query. Since Wikipedia is a common pre-training corpus for LLMs, most information might have already been encoded in the LLM. Nevertheless, the search tool can still be helpful given LLMs are known to suffer from hallucinations and the knowledge encoded might not be updated in time. For more specialized domains and sources, the search tool may be crucial. We further observe that the ablation affects WTQ and TAT more than CRT and SCITAB. This might be attributed to dataset features: CRT contains many yes-no questions and SCITAB has been converted from a ternary classification dataset. Thus, chances for guessing the correct final answers are higher than in datasets with a more diverse answer distribution, such as WTQ and TAT. By evaluating our framework on instances from CRT that have answers other than yes/no, we find a performance drop of 8.23 when ablating both tools and the coding agent.
Figure 3: EM (line chart) and iteration ratio (bar chart) against different $\alpha$. The iteration ratio is calculated by dividing the number of iterations when using efficiency optimization with a specific $\alpha$ by the number of iterations when not using it (without shortcuts).
Analysis of Intent Distribution.
We report the distribution of the intents of the selected actions for each dataset in Figure 2. We observe that Retrieve and Calculate are the most frequent intents, along with Finish. This indicates that our proposed second agent, the coding agent $M_{c}$ is used frequently. Different datasets also require different intents. In particular, the framework needs to use the intent Read for solving instances in TAT, where textual descriptions are given while this is not the case for the other datasets. We notice that the Search intent is used only few times across datasets. This might be because most instances were designed to be solved using the given table and text information. However, when looking into the individual cases where Search is used, we still find it useful. For instance, one question from WTQ asks about the number of athletes from American but no information about nationality is given in the table. In this case, Search assists with answering the question by adding the nationality for each athlete from Wikipedia. Though the intents Read, Search and Ask are less used compared to others, we still incorporate them to adapt to various use cases that might occur in real-life use cases.
@@
To investigate the trade-off between efficiency and accuracy, we plot the model performance against $\alpha\in\{0.2,0.4,0.6,0.8,1\}$ in Figure 3. We also plot the ratio of the total number of iterations taken to terminate from each tested $\alpha$ to the total number of iterations taken when not using the optimization, i.e., letting the planning agent decide via the Finish action when to stop the execution. The best performance is reached for $\alpha=1$, i.e., when requiring all estimated results to agree with each other to stop the iteration. For SCITAB, for instance, we save approximately 40% of the iterations when setting $\alpha$ to 1 without losing performance compared to not using the optimization component (59.8% vs. 59.7%). On average, adding the efficiency optimization module saves up to 33% of iterations. This shows the effectiveness of the optimization and that users can individually tune the desired trade-off of performance and computation time.
Error Analysis.
We randomly sample 50 instances that MACT fails per dataset and conduct an error analysis. About half of the errors come from invalid or wrong code generated by the coding agent $M_{c}$. Either $M_{c}$ fails to make sense of instructions or of complex table structure. The second error type can be attributed to evaluation. We find that about one-third of failures come from strict evaluation metrics (EM). This influences the performance of MACT particularly on the TAT dataset, as it features long text strings as answers. The evaluation challenge has been discussed in many previous works Wu and Feng (2024); Li et al. (2024). To estimate the upper-bound of our method, we use GPT-4 as evaluator to determine if a predicted answer is semantically the same as the reference answer. This results in an accuracy of 87.8% on the TAT dataset, compared to 66.2% with EM. The remaining error cases can be largely attributed to the failure of the planning agent in decomposing questions correctly. For instance, one question asks for the score range (min-max) of the top 10 finishers. Apart from retrieving the min and max scores of the top 10 finishers, the planner continues to generate the action: Calculate[Calculate the range of the scores in the observation 1.]. This leads to a wrong prediction.
@@ 7 Conclusions
We have proposed MACT, a multi-agent collaboration with tool use for table question answering. Unlike previous work, MACT neither requires fine-tuning nor does it depend on closed-source models. In our experiments, our framework demonstrates good generalizability across different benchmark datasets and outperforms a number of state-of-the-art approaches, including closed-source commercial models and fine-tuned models. To boost efficiency, we introduce an efficiency optimization module that saves up to 33% of the iterations in our analysis. In our experiments and analyses, we show that multi-agent collaboration with tools is an effective approach for table question answering.
8 Limitations
@@
Acknowledgements
This work was partially supported by the EU Project SMARTY (GA 101140087).
@@ A.1 Datasets
Table A.1 shows their statistics and characteristics. WTQ Pasupat and Liang (2015), TAT (Zhu et al., 2021), CRT Zhang et al. (2023b) and SCITAB (Lu et al., 2023) are publicly available under the licenses of CC-BY-SA-4.05, MIT and MIT, respectively. These licenses all permit us to compose, modify, publish, and distribute additional annotations upon the original dataset.
@@ A.3 Analysis of Iteration Number Distribution
A.5 Case Study
We present two reasoning traces selected from CRT and WTQ in Figure 4 and 5, respectively. Figure 4 presents a case involving complex reasoning, where models need to identify top ten finishers, calculate percentage of drivers for each constructor and return the constructor with the highest percentage. In Figure 5, we observe the planning agent has the ability to self-correct in Thought 3.
Figure 4: An instance selected from CRT featuring complex reasoning.
Figure 5: An instance selected from WTQ. We find the planning agent can perform self-correct given previous reasoning traces.
@@ A.6 MACT Prompts
We provide the prompts used for the planning agent for the examined datasets (WTQ, TAT, CRT and SCITAB) in Figures 8, 9, 10, and 11. Figure 12 and Figure 13 show the prompts used for the coding agent for the action intents Retrieval and Calculation, respectively.